# Self-Supervised Frameworks for Speaker Verification via Bootstrapped Positive Sampling

### v2

@@ V-B1 SSPS-NN
SSPS-NN samples a pseudo-positive close to the anchor in the latent space with different sampling window sizes $M$. The best configuration is obtained with $M=50$, which implies that the algorithm should sample positives close to the anchor to ensure that they belong to the same speaker identity but still have a large sampling window to find positives from different recording sources. There is a clear trade-off between speaker and recording accuracy, as having a small sampling window can prevent the method from selecting different recordings, which is highlighted by Figure 4a. With $M=50$, SSPS-NN effectively improves SV performance and provides pseudo-positives with a speaker accuracy of around 50% and a recording accuracy of around 25%.
@@ V-B2 SSPS-Clustering
SSPS-Clustering samples a pseudo-positive according to the clustering assignments of the latent representations. Using $K=6{,}000$, which approximately corresponds to the number of speakers in the train set, and sampling from the same cluster as the anchor with $M=0$, the EER is reduced to 6.63%. Setting $K=25{,}000$ and $M=1$ further improves the EER to 5.80% and corresponds to the best configuration for the minDCF. However, setting $K=150{,}000$, which represents the actual number of recordings in the train set, results in worse performance, which may be caused by the fact that several recordings may be included in the same cluster. As shown in Figure 4b, using a large value of $K$ and sampling from a neighboring cluster, i.e., $M>0$, the overall performance is better, and the pseudo-positives speaker accuracy is maintained while the recording accuracy drops compared to SSPS-NN. Note that using the sampling cluster centroid, i.e., SSPS-Clustering (C), as a positive embedding is ineffective, suggesting that the embedding space is dense but not continuous.
In the following, only SSPS-Clustering will be applied, as it allows sampling pseudo-positive with higher speaker accuracy and lower recording accuracy compared to SSPS-NN.
@@ V-C Evaluation of SSPS on SV
TABLE III: Evaluation of SSL frameworks on SV with different positive sampling strategies: SSL, SSPS, and Supervised. The encoder is ECAPA-TDNN, and the benchmarks are VoxCeleb1-O, VoxCeleb1-E, and VoxCeleb1-H. Best results are in bold, excluding supervised baselines.
Method Positive sampling SSPS Hyper-params. VoxCeleb1-O VoxCeleb1-E VoxCeleb1-H
@@
In Table III, the performance of the main SSL frameworks for different configurations of the SSPS-Clustering strategy is reported on SV using the ECAPA-TDNN encoder.
Configurations using $K=25{,}000$ and $M=1$ outperform those using $K=6{,}000$ and $M=0$, which confirms the assumption that speaker representations are grouped according to channel information in latent space. In the following, SSPS-Clustering is referred to as SSPS with the former set of hyperparameters, as it corresponds to the best performance on the downstream task.
Compared to SSL positive sampling, which corresponds to the baseline, SSPS improves the EER and minDCF of all reported SSL methods on all VoxCeleb benchmarks. This shows the effectiveness of the proposed approach when applied to various SSL frameworks. As expected, the best result is obtained by DINO with SSPS, achieving 2.53% EER and 0.2843% minDCF on VoxCeleb1-O. Note that this performance is very close to its Supervised positive sampling baseline, resulting in 2.36% EER and 0.2712 minDCF. Furthermore, SimCLR provides comparable performance to DINO by reaching 2.57% EER and 0.3033 minDCF. The performance of SwAV and VICReg are also improved with SSPS, from 7.97% to 6.50% EER and from 7.70% to 6.95% EER, respectively. The lower performance improvement on SwAV and VICReg can be attributed to the already limited performance of their Supervised positive sampling baselines, suggesting that the same-utterance positive sampling is not the only bottleneck. Therefore, SSPS proves to be very effective in improving the downstream performance of major SSL frameworks on SV. For comparison, the supervised method achieves 1.34% EER and 0.1521 minDCF on VoxCeleb1-O.
SimCLR, based on the contrastive learning paradigm, demonstrates a remarkable improvement over its baseline when using SSPS, as it results in a 58% relative EER reduction. This is very promising as this system outperforms DINO without SSPS, even though DINO relies on a far more complex training framework. The Supervised positive sampling strategy applied to SimCLR achieves 1.72% EER and 0.2395 minDCF, implying that there is potential for further improvements of SSL contrastive-based methods.
Finally, it can be observed that: (1) DINO with SSPS reaches a performance comparable to its Supervised positive sampling baseline, suggesting that this approach has reached its full potential; (2) SimCLR with Supervised positive sampling results in the best performance among SSL methods, outperforming DINO with Supervised positive sampling. This shows that the focus should be on contrastive SSL frameworks for SV, as these techniques benefit from the discriminative capacity of the contrastive-based objective function. Thus, the following experiments to demonstrate the effect of SSPS will be conducted on SimCLR using the ECAPA-TDNN encoder.
@@ V-D Comparison to state-of-the-art SSL methods for SV
TABLE IV: Comparison of state-of-the-art SSL methods for SV. Methods are grouped into contrastive and self-distillation frameworks. The benchmark is VoxCeleb1-O. Results drawn from the literature are highlighted in gray and the proposed methods are in bold.
@@
AP + AAT [huh2020APAAT] $8.65$
Contrastive + VICReg [lepage2022LabelEfficientSSLSV] $8.47$ $0.6400$
SimCLR + MSE loss [zhang2021SimCLR] $8.28$ $0.6100$
MoCo + ProtoNCE [xia2021MoCo] $8.23$ $0.5900$
CEL [mun2020CEL] $8.01$
SimCLR + AAT [tao2022LGL] $7.36$
C3-MoCo [zhang2022C3DINO] $6.40$
DPP [tao2023DPP] $2.89$
@@
SSReg [sang2022SSReg] $6.99$
DINO + Cosine loss [han2022DLGLC] $6.16$ $0.5240$
DINO [cho2022DINO] $4.83$ $0.4630$
DINO + Curriculum [heo2022DINOCurriculum] $4.47$
CA-DINO [han2024CADINO] $3.59$ $0.3529$
RDINO [chen2023RDINO] $3.29$
MeMo [jin2024WGVKT] $3.10$
RDINO + W-GVKT [jin2024MeMo] $2.89$ $0.3330$
DINO + RMP [kim2024RMPSSLSV] $2.62$
@@
DINO + Aug. [chen2022ComprehensiveStudySelfDistillation] $2.51$
C3-DINO [zhang2022C3DINO] $2.50$
@@
Among contrastive methods, the proposed SimCLR-SSPS outperforms all prior approaches and sets a new state-of-the-art of 2.57% EER on the benchmark. This confirms the effectiveness of the bootstrapped positive sampling strategy within SSL contrastive frameworks. Although DPP [tao2023DPP] achieves a comparable EER of 2.89%, its positive sampling leverages audio-visual cues from speech-face pairs of VoxCeleb2, whereas SSPS is trained using audio data only.
In the self-distillation category, the proposed DINO-SSPS also demonstrate competitive performance by reaching 2.53% EER compared to the other methods listed. While DINO + Aug. [chen2022ComprehensiveStudySelfDistillation] and C3-DINO [zhang2022C3DINO] achieve marginally lower EERs (2.51% and 2.50%, respectively), their performance is comparable to DINO-SSPS given the minimal difference. Notably, the former depends on a more complex data-augmentation strategy, and the latter initializes the DINO model from pre-trained MoCo weights, whereas DINO-SSPS attains competitive results without these specific enhancements.
@@ V-E Effect of SSPS on SSL speaker representations
The median cosine similarity of same-speaker representations for train and test samples increases when using SSPS compared to SSL positive sampling, as shown in Figure 5. This effect was expected as providing more diverse positives to the SSL framework allows matching different representations with distinct channel characteristics to the same speaker identity and thus reduces the intra-class variability. This improvement in class compactness is visible in Figure 6, which represents the t-SNE of 10 speakers representations from VoxCeleb1 with SSL and SSPS positive sampling techniques. Some groups of samples (“sub-clusters”) remain distant in the latent space despite belonging to the same speaker identity, suggesting that SSPS may not effectively consider all utterances from a given speaker as pseudo-positives. This demonstrates a limitation of the proposed approach whenever the SSL model has already separated same-speaker representations in latent space, due to strong acoustic variability, prior to the activation of SSPS.
@@
Figure 7: NMI ratio of speaker and recording information with SSL and SSPS across training epochs. The model is SimCLR, the encoder is ECAPA-TDNN, and the NMI is computed with $K=6{,}000$ clusters on reference representations extracted from VoxCeleb2.
To further assess the quality of the SSL speaker representations learned with SSPS, the Normalized Mutual Information (NMI) [strehl2002ClusterEnsembles] between the clustering assignments of the reference representations, obtained with $K=6{,}000$, and the VoxCeleb2 metadata, including speaker and recording labels, is used to quantify the alignment with the underlying structure of the data. The NMI between two sets of assignments $U$ and $V$ is defined as:
@@
where $I(U;V)$ is the Mutual Information (MI) between $U$ and $V$, $H(U)$ is the entropy of $U$, and $H(V)$ is the entropy of $V$. The NMI should be maximized, as it reflects the degree of alignment between the predicted cluster assignments and the ground-truth labels. This allows reporting the ratio of speaker-to-recording NMI across epochs with SSL and SSPS positive sampling strategies in Figure 7. Before the activation of SSPS, the NMI ratio of $\sim$1.01 implies that the representations contain as much information on speaker identity as on the recording source. At the end of the training, SSPS has increased the ratio of speaker over recording NMI to $\sim$1.09, which represents a $\sim$8relative improvement, while the SSL baseline results in a ratio close to the initial value. Thus, SSPS allows the model to learn representations that are more robust to extrinsic variabilities arising from the different recording conditions by encoding less channel information compared to the SSL same-utterance positive sampling.
@@ V-F Robustness of SSPS to the absence of data-augmentation
This experiment evaluates the ability of the proposed approach to converge without any data-augmentation pre-processing steps by comparing SSL and SSPS positive sampling techniques with their counterparts without data-augmentation. The results reported with SimCLR (ECAPA-TDNN) in Table V show that, without data-augmentation, the performance of the baseline significantly degrades from 6.30% to 15.00% EER on VoxCeleb1-O, whereas SSPS maintains strong performance with only a minor degradation of the EER from 2.57% to 2.77%, demonstrating its robustness to the absence of data-augmentation. Notably, the minDCF of SSPS even improves slightly when no data-augmentation is employed, decreasing from 0.3033 to 0.2840. This highlights the effectiveness of SSPS as it outperforms the same-utterance positive sampling used in SSL, while simultaneously eliminating the need to artificially create diverse acoustic conditions.
The EER on VoxCeleb1-O across epochs for these systems is shown in Figure 8. The model based on SSL positive sampling without data-augmentation (light gray, dashed) suffers from the similarity of channel characteristics between the anchor and the positive, leading to performance degradation from the first epoch. In contrast, the model based on SSPS maintains consistent performance, with (blue, solid) and without (light blue, dashed) data-augmentation, throughout training. This figure underscores the strong reliance of the standard same-utterance positive sampling on data-augmentation and justifies the underlying intuition of SSPS to create anchor-positive pairs from different recordings of the same speaker instead of relying on standard augmentation techniques.
The reduced reliance of the proposed method on data-augmentation is an important property, considering that data-augmentation is a fundamental component for training SSL frameworks. However, data-augmentation has several drawbacks: it depends heavily on the downstream task and input data, and often involves a hyperparameter space too large to be explored efficiently in many scenarios. This suggests that alternative combinations of positive sampling and data-augmentation strategies have to be developed.
@@ VI Conclusions
To overcome the limitations of the default SSL same-utterance positive sampling, Self-Supervised Positive Sampling (SSPS) is proposed as a new strategy for sampling positives in SSL frameworks. For SV, this new approach finds relevant and diverse pseudo-positives in the latent space that are not derived from the same recording as the anchor, which reduces the channel and recording information in speaker representations. SSPS improves downstream performance for all major SSL methods, i.e., SimCLR, SwAV, VICReg, and DINO, on VoxCeleb benchmarks compared to the default positive sampling. The best performance is achieved with SimCLR and DINO, reaching, respectively, 2.57% and 2.53% EER on VoxCeleb1-O, where the supervised baseline reaches 1.34% EER. In particular, SimCLR with SSPS results in a 58% relative reduction in EER. When using Supervised positive sampling, better performance is achieved with SimCLR compared to DINO, significantly reducing the gap with the fully supervised baseline, which motivates the need to consider contrastive-based SSL frameworks for the task of SV. Furthermore, better class compactness and a reduction of channel and recording information from the speaker representations are observed. Finally, the proposed method is shown to be more robust to the absence of data-augmentation than standard SSL models. This work addresses key limitations of SSL methods for SV, advancing the field toward closing the performance gap between supervised and self-supervised systems.

### v3

@@ V-B1 SSPS-NN
SSPS-NN samples a pseudo-positive close to the anchor in the latent space with different sampling window sizes $M$. The optimal configuration is achieved with $M=50$, which implies that the algorithm should sample pseudo-positives close to the anchor to ensure they belong to the same speaker identity, while still maintaining a large sampling window to identify pseudo-positives from different recording sources. There is a clear trade-off between pseudo-positive speaker and pseudo-positive recording accuracies, as a small sampling window can prevent the method from selecting different recordings, as highlighted in Figure LABEL:sub@fig:ssps_nn_hyperparams. With $M=50$, SSPS-NN effectively improves SV performance, providing pseudo-positives with a speaker accuracy of approximately 50% and a recording accuracy of approximately 25%.
@@ V-B2 SSPS-Clustering
[…]
@@ V-C Evaluation of SSPS on SV
TABLE III: Evaluation of SSL frameworks on SV with different positive sampling strategies: SSL, SSPS, and Supervised. The encoder is ECAPA-TDNN, and the benchmarks are VoxCeleb1-O, VoxCeleb1-E, and VoxCeleb1-H. The best methods are highlighted in light blue and the baselines are highlighted in light gray. Best results are in bold, excluding supervised baselines.
Framework Pos. sampling SSPS Hyperparameters VoxCeleb1-O VoxCeleb1-E VoxCeleb1-H
@@
The performance of SSL frameworks on SV with different configurations of the SSPS-Clustering strategy is reported using the ECAPA-TDNN encoder in Table III.
[…]
Compared to SSL positive sampling, which serves as the baseline, SSPS improves the EER and minDCF of all reported SSL frameworks across all VoxCeleb benchmarks. This shows the effectiveness of the proposed approach when applied to various SSL paradigms. As expected, the best result is achieved by DINO with SSPS, yielding 2.53% EER and 0.2843% minDCF on VoxCeleb1-O. Note that this performance is very close to its Supervised positive sampling baseline, resulting in 2.36% EER and 0.2712 minDCF. Furthermore, SimCLR provides comparable performance to DINO by reaching 2.57% EER and 0.3033 minDCF. The performance of SwAV and VICReg is also improved with SSPS, from 7.97% to 6.50% EER and from 7.70% to 6.95% EER, respectively. The lower performance improvement on SwAV and VICReg can be attributed to the already limited performance of their Supervised positive sampling baselines, suggesting that the same-utterance positive sampling is not the only bottleneck. Therefore, SSPS proves to be very effective in improving the downstream performance of major SSL frameworks on SV. For comparison, the supervised method achieves 1.34% EER and 0.1521 minDCF on VoxCeleb1-O.
SimCLR, based on the contrastive learning paradigm, demonstrates a remarkable improvement over its baseline when using SSPS, as it results in a 58% relative EER reduction. This is very promising, as the system outperforms DINO without SSPS, despite DINO relying on a far more complex training framework. Supervised positive sampling strategy applied to SimCLR achieves 1.72% EER and 0.2395 minDCF, implying that there is potential for further improvements of SSL contrastive-based methods.
Finally, it can be observed that: (1) DINO with SSPS reaches a performance comparable to its Supervised positive sampling baseline, suggesting that this approach has reached its full potential; (2) SimCLR with Supervised positive sampling results in the best performance among SSL methods, outperforming DINO with Supervised positive sampling. This shows that the focus should be on SSL contrastive frameworks, which benefit from the discriminative capacity of the contrastive-based objective function. Therefore, the following experiments will be conducted with the SimCLR framework.
@@ V-D Comparison to state-of-the-art SSL methods for SV
TABLE IV: Comparison of state-of-the-art SSL methods for SV. Methods are grouped into contrastive and self-distillation frameworks. The benchmark is VoxCeleb1-O, with EERs sorted in descending order. The proposed methods are in bold and highlighted in light blue. Other results are drawn from the literature.
@@
AP + AAT [14] $8.65$
Contrastive + VICReg [35] $8.47$ $0.6400$
SimCLR + MSE loss [15] $8.28$ $0.6100$
MoCo + ProtoNCE [16] $8.23$ $0.5900$
CEL [39] $8.01$
SimCLR + AAT [40] $7.36$
C3-MoCo [18] $6.40$
DPP [24] $2.89$
@@
SSReg [41] $6.99$
DINO + Cosine loss [42] $6.16$ $0.5240$
DINO [21] $4.83$ $0.4630$
DINO + Curriculum [22] $4.47$
CA-DINO [20] $3.59$ $0.3529$
RDINO [19] $3.29$
MeMo [43] $3.10$
RDINO + W-GVKT [44] $2.89$ $0.3330$
DINO + RMP [45] $2.62$
@@
DINO + Aug. [46] $2.51$
C3-DINO [18] $2.50$
@@
Among contrastive methods, the proposed SimCLR-SSPS outperforms all prior approaches and sets a new state-of-the-art of 2.57% EER on the benchmark. This confirms the effectiveness of the bootstrapped positive sampling strategy within SSL contrastive frameworks. Although DPP [24] achieves a comparable EER of 2.89%, its positive sampling leverages audio-visual cues from speech-face pairs of VoxCeleb2, whereas SSPS is trained using audio data only.
In the self-distillation category, the proposed DINO-SSPS also demonstrates competitive performance compared to the other methods listed by reaching 2.53% EER. While DINO + Aug. [46] and C3-DINO [18] achieve marginally lower EERs (2.51% and 2.50%, respectively), their performance is comparable to DINO-SSPS, given the minimal difference. Notably, the former relies on a more complex data-augmentation strategy, while the latter initializes the DINO model from pre-trained MoCo weights. In contrast, DINO-SSPS achieves competitive results without these specific enhancements.
@@ V-E Effect of SSPS on SSL speaker representations
The median cosine similarity of same-speaker representations for train and test samples increases when using SSPS compared to SSL positive sampling, as shown in Figure 5. This effect was expected, as providing more diverse positives to the SSL framework allows matching different representations with distinct channel characteristics to the same speaker identity, thereby reducing intra-class variability. This improvement in class compactness is visible in Figure 6, which represents the t-SNE of 10 speakers’ representations from VoxCeleb1 with SSL and SSPS positive sampling techniques. Some groups of samples (“sub-clusters”) remain distant in the latent space despite belonging to the same speaker identity, suggesting that SSPS may not effectively consider all utterances from a given speaker as pseudo-positives. This demonstrates a limitation of the proposed approach whenever the SSL model has already separated same-speaker representations in the latent space, due to strong acoustic variability, prior to the activation of SSPS.
@@
Figure 7: NMI ratio of speaker and recording information with SSL and SSPS across training epochs. The model is SimCLR, the encoder is ECAPA-TDNN, and the NMI is computed with $K=$6000$$ clusters on reference representations extracted from VoxCeleb2.
To further assess the quality of the speaker representations learned with SSPS, the Normalized Mutual Information (NMI) [47] between the clustering assignments of the reference representations, obtained with $K=$6000$$, and the VoxCeleb2 metadata, including speaker and recording labels, is used to quantify the alignment with the underlying structure of the data. The NMI between two sets of assignments $U$ and $V$ is defined as:
@@
where $I(U;V)$ is the Mutual Information (MI) between $U$ and $V$, $H(U)$ is the entropy of $U$, and $H(V)$ is the entropy of $V$. The NMI should be maximized, as it reflects the degree of alignment between the predicted cluster assignments and the ground-truth labels. This allows reporting the ratio of speaker-to-recording NMI across epochs with SSL and SSPS positive sampling strategies in Figure 7. Before the activation of SSPS, the NMI ratio of $\sim$1.01 implies that the representations contain as much information on speaker identity as on the recording source. At the end of the training, SSPS has increased the ratio of speaker over recording NMI to $\sim$1.09, which represents a $\sim$8relative improvement, while the SSL baseline results in a ratio close to the initial value. Thus, SSPS enables the model to learn representations that are more robust to extrinsic variabilities arising from different recording conditions by encoding less channel information compared to SSL same-utterance positive sampling.
@@ V-F Robustness of SSPS to the absence of data-augmentation
This experiment evaluates the ability of the proposed approach to converge without any data-augmentation pre-processing steps by comparing SSL and SSPS positive sampling techniques with their counterparts that do not involve data-augmentation. The results reported with SimCLR (ECAPA-TDNN) in Table V show that, without data-augmentation, the performance of the baseline significantly degrades from 6.30% to 15.00% EER on VoxCeleb1-O, whereas SSPS maintains strong performance with only a minor degradation of the EER from 2.57% to 2.77%, demonstrating its robustness to the absence of data-augmentation. Notably, the minDCF of SSPS even improves slightly when no data-augmentation is employed, decreasing from 0.3033 to 0.2840. This highlights the effectiveness of SSPS, as it outperforms the same-utterance positive sampling used in SSL, while simultaneously eliminating the need to artificially create diverse acoustic conditions.
The EER on VoxCeleb1-O across epochs for these systems is shown in Figure 8. The model based on SSL positive sampling without data-augmentation (light gray, dashed) suffers from the similarity of channel characteristics between the anchor and the positive, leading to performance degradation from the first epoch. In contrast, the model based on SSPS maintains consistent performance, with (blue, solid) and without (light blue, dashed) data-augmentation, throughout training. This figure highlights the strong reliance of the standard same-utterance positive sampling on data-augmentation. This justifies the underlying intuition of SSPS, which is to create anchor-positive pairs from different recordings of the same speaker, rather than relying on augmentation techniques.
The reduced reliance of the proposed method on data-augmentation is an important property, considering that data-augmentation is a fundamental component for training SSL frameworks. However, data-augmentation is not an ideal approach as it depends heavily on the downstream task and input data, and often involves a hyperparameter space too large to be explored efficiently in many scenarios. This suggests that alternative combinations of positive sampling and data-augmentation strategies should be investigated.
@@ VI Conclusions
To overcome the limitations of the default SSL same-utterance positive sampling, Self-Supervised Positive Sampling (SSPS) is proposed as a new strategy for sampling positives in SSL frameworks. For SV, this new approach identifies relevant and diverse pseudo-positives in the latent space that are not derived from the same recording as the anchor, which reduces the channel and recording information in speaker representations. SSPS enhances downstream performance for all major SSL frameworks, including SimCLR, SwAV, VICReg, and DINO, on VoxCeleb benchmarks compared to the default positive sampling approach. The best performance is achieved with SimCLR and DINO, reaching 2.57% and 2.53% EER, respectively, on VoxCeleb1-O, where the supervised baseline achieves 1.34% EER. Notably, SimCLR with SSPS yields a 58% relative reduction in EER. When using Supervised positive sampling, better performance is achieved with SimCLR compared to DINO, significantly reducing the gap with the fully supervised baseline, which motivates the need to consider contrastive-based SSL frameworks for the task of SV. Furthermore, better class compactness and a reduction of channel and recording information from the speaker representations are observed. Finally, the proposed method is shown to be more robust to the absence of data-augmentation than standard SSL models. This work addresses key limitations of SSL methods for SV, advancing the field toward closing the performance gap between supervised and self-supervised systems.